# Mitigating the Impact of Labeling Errors on Training via Rockafellian Relaxation

### v1

@@ 6.1 MNIST
Table 1: Test accuracy (%) for AT and A-RRM on MNIST under different levels of corruption $C$ and test-set adversarial perturbation $\epsilon_{test}$.
Percentage Corrupted Training Data
$C$ 0% 5% 10% 20% 30%
$\epsilon_{test}$ AT A-RRM AT A-RRM AT A-RRM AT A-RRM AT A-RRM
0.00 97 96 63 95 57 97 58 96 26 86
@@
We examine the $u_{i}$-value associated with each training observation, $i$, from iteration-to-iteration of the heuristic algorithm. Table 2 summarizes the progression of the $u_{i}$-vector across its 49 updates for the dataset corruption level of 20%. Column “1. iteration” shows the distribution of $u_{i}$-values following the first u-optimization for both the 9600 corrupted training observations and the 38400 clean training observations. Initially, all $u_{i}$-values are approximately equal to 0.0. It is once again observed that, over the course of iterations, the $u_{i}$-values noticeably change. In column “10. iteration” it can be seen that a significant number of the $u_{i}$-values of the corrupted training observations achieve negative values, while a large majority of the $u_{i}$-values for the clean training observations remain close to 0.0. Finally, column “49. iteration” displays the final $u_{i}$-values. 9286 out of 9600 of the corrupted training observations have achieved a $u_{i}\in(-2.08,-1.56]\cdot 10-5$. This means these training observations are removed, or nearly-so, from consideration because this value cancels the nominal probability $1/N$ = 2.08 $\cdot$ 10-5. It is observed that a large majority (35246/38400) clean training observations remain with their nominal probability. This helps explain the performance benefit of A-RRM over AT. A-RRM "removes" the corrupted data points in-situ, whereas AT does not. It appears that under adversarial training regimes with corrupted training data, it is essential to identify and "remove" the corrupted examples, especially if the level adversarial perturbation encountered in the test set is unknown, or possibly lower than the level of adversarial perturbation applied to the training set.
Table 2: Evolution of u-vector across 9600 corrupted data points and 38400 clean data points. Note that 1/(9600 + 38400) $=$ 2.08 $\cdot$ 10-5.
1. iteration 10. iteration 49. iteration
$u_{i}$ value corrupted data points clean data points corrupted data points clean data points corrupted data points clean data points
$\gg$ 0 0 1 0 4. 0 25
$\approx$ 0 8844 38385 2058 37524 91 35246
(-0.52, 0.00) $\cdot$ 10-5 0 0 7 36 146 1655
(-1.04, -0.52] $\cdot$ 10-5 0 0 41 45 43 155
(-1.56, -1.04] $\cdot$ 10-5 756 14 415 174 34 168
(-2.08, -1.56] $\cdot$ 10-5 0 0 7079 617 9286 1151
6.2 Toxic Comment
We use the Toxic Comment dataset to test the efficacy of RRM on low prevalence text data. The positive (toxic) comments consist of only 3% of the data and we corrupt anywhere from 1% to 20% of the labels. There are a total of 148,000 samples, and we set aside 80% for training and 20% for test. $\sigma=2$ with 3 iterations of the heuristic algorithm results in a total of 6 epochs, and ERM is run for a total of 6 epochs to make the results comparable. Since the data is highly imbalanced, we look at the area under the curve of the precision/recall curve to assess the performance of the models. Unsurprisingly, as the noise increase, the model performance decreases. We note that RRM outperforms ERM across all noise levels tested, though as the noise increase, the gap between RRM and ERM decreases.
Table 3: Comparison of training and test area under the precision/recall curve for ERM and RRM at noise levels ranging from 1% to 20%.
Method Percentage Corrupted Training Data
1% 5% 7% 10% 15% 20%
ERM (train) 0.2904 0.2006 0.1589 0.1302 0.1073 0.0920
RRM (train) 0.6875 0.4458 0.3805 0.3087 0.2438 0.1966
ERM (test) 0.5861 0.3970 0.3246 0.2550 0.2013 0.1717
RRM (test) 0.6705 0.4338 0.3619 0.2824 0.2208 0.1861
6.3 IMDb
Twenty percent of the training data is set aside for validation purposes. Using Pytorch 2.1.0 [3], 30 iterations of RRM are executed, with $\sigma=10$ epochs per iteration for a total of 300 epochs for a given hyperparameter setting. For RRM, the hyperparameter settings of $\mu$ and $\gamma$ at 0.5 and 0.4, respectively, are based on a search to optimize validation set accuracy. For contrast, we perform a comparable 300 epochs using ERM. Both ERM and RRM employ stochastic gradient descent (SGD) with a learning rate ($\eta$) of $0.001$. In Table 4 we record both the test set accuracy achieved when validation set accuracy peaks, as well as the maximum test set accuracy. At these high levels of corruption RRM consistently achieves a better maximum test set accuracy.
Table 4: Test accuracy (%) for ERM and RRM on IMDb under different levels of corruption. Test set accuracy at peak validation accuracy and maximum test set accuracy are recorded.
Method Percentage Corrupted Training Data
25% 30% 40% 45%
ERM 90.2, 90.2 89.5, 89.6 86.4, 86.6 80.7, 81.1
RRM 90.1, 90.4 90.2, 90.4 88.4, 88.7 76.9, 82.6
6.4 Tissue Necrosis
@@
Twenty percent of the training data is set aside for validation purposes, including hyperparameter selection. 60 iterations of RRM are executed, with $\sigma=10$ epochs per iteration, for a total of 600 epochs for a given hyperparameter setting. For RRM, the hyperparameter settings of $\mu$ and $\gamma$ at 0.5 and 0.016, respectively, are based on a search to optimize validation set accuracy. For contrast, we perform a comparable 600 epochs using ERM. Both ERM and RRM employ stochastic gradient descent (SGD) with a learning rate ($\eta$) of $5.0$ and $1.0$, respectively. RRM achieves a test set accuracy at peak validation accuracy of 74.6, and a maximum test set accuracy 77.2, whereas ERM achieves 71.7 and 73.2, respectively. RRM appears to confer a performance benefit under this regime of weakly labeled data.
@@ 7 Conclusion
In this study, we demonstrate the robustness of the A-RRM algorithm in a variety of data domains, data corruption schemes, model architectures and machine learning applications. In the MNIST example we show that conducting training in preparation for deployment environments with varied levels of adversarial attacks, one can benefit from implementation of the A-RRM algorithm. This can lead to a model more robust across levels of both feature perturbation and high levels of label corruption. We also demonstrate the mechanism by which A-RRM operates and confers superior results: by automatically identifying and removing the corrupted training observations at training time execution.
The Toxic Comment example presents another challenging classification problem, characterized by a low prevalence target class amidst label noise. Our experiments demonstrate that as the amount of label noise increases, standard methods become increasingly ineffective. However, RRM remains reasonably robust under varying degrees of label corruption. Therefore, RRM could be a valuable addition to the set of tools being developed to enhance the robustness of AI-based decision engines.
In the IMDb example we demonstrate that RRM can confer benefits to the sentiment analysis classification task using pre-trained large models under conditions of high label corruption. The success of fine-tuning in LLMs depends, in large part, on access to high quality training examples. We have shown that RRM can mitigate this need by allowing effective training in scenarios of high training data corruption. As such, resource allocation dedicated to dataset curation may be lessened by the usage of RRM.
In the Tissue Necrosis example, we demonstrate that RRM also confers accuracy benefits to the necrosis identification task provided weakly labeled WSIs. Again, RRM can mitigate the need for expert-curated, detailed pathology annotations, which are costly and time-consuming to generate.
@@ A.1 Section 3 Proofs
$h(u)=\sum_{i=1}^{N}\Big{(}c_{i}/N+u_{i}c_{i}+\gamma|u_{i}|+\iota_{[0,\infty)}(1/N+u_{i})\Big{)}+\iota_{\{0\}}\Big{(}\sum_{i=1}^{N}u_{i}\Big{)}$
Thus, because $h(u)>-\infty$ for all $u\in\mathbb{R}^{N}$ and $h$ is convex, $u^{\star}$ is a solution of the minimization problem if and only if $0\in\partial h(u^{\star})$ by Theorem 2.19 in [24]. We proceed by characterizing $\partial h$.
@@
$h_{i}(u_{i})=c_{i}/N+u_{i}c_{i}+\gamma|u_{i}|+\iota_{[0,\infty)}(1/N+u_{i}).$
For $u_{i}\geq-1/N$, the Moreau-Rockafellar sum rule (see, e.g, [24, Theorem 2.26]) gives that
$\partial h_{i}(u_{i})=c_{i}+\begin{cases}\{\gamma\}&\mbox{ if }u_{i}>0\\ [-\gamma,\gamma]&\mbox{ if }u_{i}=0\\ \{-\gamma\}&\mbox{ if }-1/N<u_{i}<0\\ (-\infty,-\gamma]&\mbox{ if }u_{i}=-1/N.\end{cases}$
For $u=(u_{1},\dots,u_{N})\in[-1/N,\infty)^{N}$, we obtain by Proposition 4.63 in [24] that
@@
$\lambda\in\begin{cases}\{c_{i}+\gamma\}&\mbox{ if }u_{i}^{\star}>0\\ [c_{i}-\gamma,c_{i}+\gamma]&\mbox{ if }u_{i}^{\star}=0\\ \{c_{i}-\gamma\}&\mbox{ if }u_{i}^{\star}\in(-1/N,0)\\ (-\infty,c_{i}-\gamma]&\mbox{ if }u_{i}^{\star}=-1/N.\end{cases}$
It follows that $\lambda=c_{min}+\gamma$ can accompany any optimal $u^{*}$ in satisfying the above; hence, the result follows.
See 3.2
@@
By Lemma 3.1 of [30] and/or Corollary 2 of [9],

### v2

@@ 6.2 Enhancing Loss Function Methodologies with RRM
Table 4: MNIST-10 Test accuracy (%) for AT and A-RRM under different levels of contamination $C$, training perturbation $\epsilon=1.0$, and test-set adversarial perturbation $\epsilon_{test}$.
$C$
0% 5% 10% 20% 30%
AT A-RRM AT A-RRM AT A-RRM AT A-RRM AT A-RRM
$\epsilon_{test}$ 0.00 97 96 63 95 57 97 58 96 26 86
@@
Implementation Details: In this experiment, we divide MNIST-3 into a training set of 18623 examples (5923 0’s, 6742 1’s, and 5958 2’s), and 3147 testing examples. Using Tensorflow 2.10 (Abadi et al., 2015), 100 epochs of training are performed for the MNIST-3 architecture outlined in Section 5 for each of the baseline loss functions (ELR, CCE, MAE, and MSE) and training label contamination levels of 55%, 60%, and 65% for a total 12 baseline experiments. In order the to assess the benefit of our method, the RRM algorithm is executed for each of the baseline loss functions and contamination levels. More specifically, 10 iterations of RRM are executed with $\sigma=10$ epochs per iteration for a total of 100 epochs. For RRM, the hyperparameter settings of $\mu$ and $\gamma$ are set to 0.5 and 0.4, although auto-tuning could have been implemented as in Section 6.1. All methods employ stochastic gradient descent (SGD) with a learning rate ($\eta$) of $0.1$. The test set accuracy of each experiment is shown in Table 3, with the RRM experimental results placed in parentheses.
Analysis: For every method tested, the RRM-wrapped approach confers a test set performance benefit over the baseline approach. The additional computation time incurred from the Re-weight($\gamma,\mu$) step is an average of 2.86 seconds per execution of every iteration, for a total of 28.6 additional seconds.
6.3 A-RRM: Granting ERM Robustness to Adversarial and Noisy Data Regime
MNIST-10 is utilized to evaluate our method in a setting of both adversarial feature perturbation and label contamination.
Adversarial Training (AT) Benchmark: As benchmark, we execute a standard adversarial training (AT) approach, in which we execute A-RRM, omitting the Re-weight step. Indeed, it is common to perform training on data augmented with adversarial examples crafted through FGSM (Madry et al., 2018). Further, using this benchmark, we may highlight the value of re-weighting.
Test-set Perturbation $\epsilon_{test}$: In the experiments, upon conclusion of the A-RRM training for which a trained $\theta^{*}$ is obtained, we executed an FGSM attack on each member $(x,y)$ of the test set via $x+\epsilon_{test}\cdot\nabla J(\theta^{*};x,y)$.
In this way, in analogous fashion to the possible misalignment between estimate $C^{\prime}$ and the actual contamination level $C$, we explore the harm to performance that a misalignment between our training perturbation $\epsilon$ and a true test-set perturbation $\epsilon_{test}$ could have.
@@ 6.3 A-RRM: Granting ERM Robustness to Adversarial and Noisy Data Regime
Implementation Details: In this experiment twenty percent of the training data is set aside for validation purposes. Using Tensorflow 2.10 (Abadi et al., 2015), 50 iterations of A-RRM are executed with $\sigma=10$ epochs per iteration for a total of 500 epochs for a given hyperparameter setting. For A-RRM, the hyperparameter settings of $\mu$ and $\gamma$ at 0.5 and 2.0, respectively, are based on a search to optimize validation set accuracy. For contrast, we perform a comparable 500 epochs using AT. Both AT and A-RRM employ stochastic gradient descent (SGD) with a learning rate ($\eta$) of $0.1$. An $\epsilon=1.0$ is used for all training image perturbations.
Performance Analysis: For each of the 0%, 5%, 10%, 20%, and 30% training label contamination levels, we compare adversarial training (AT) and A-RRM performance under various regimes of test set perturbation ($\epsilon_{test}\in\{0.0,0.1,0.25,0.5,1.0\}$). In Table 4 we show the test set accuracy achieved when validation set accuracy peaks. We can see that training with an $\epsilon=1.0$ and testing with lower $\epsilon_{test}$ levels of $0.00,0.10,$ and $0.25$, results in a drastic degradation in accuracy for AT for contamination levels greater than 0%. This performance collapse is not observed when using A-RRM. Given that the $\epsilon$ employed during training may not match test/production environment ($\epsilon_{test}$), our findings suggest that A-RRM can confer a greater benefit than AT in these scenarios.
Convergence of $u$: We examine the $u_{i}$-value associated with each training observation $i$ from iteration-to-iteration of Algorithm 1. Table 5 summarizes the progression of the $u_{i}$-vector across its 49 updates for the dataset contamination level of 20%. Column “1. iteration” shows the distribution of $u_{i}$-values following the first u-optimization for both the 9600 contaminated training observations and the 38400 clean training observations. Initially, all $u_{i}$-values are approximately equal to 0.0. It is once again observed that, over the course of iterations, the $u_{i}$-values noticeably change. In column “10. iteration” it can be seen that a significant number of the $u_{i}$-values of the contaminated training observations achieve negative values, while a large majority of the $u_{i}$-values for the clean training observations remain close to 0.0. Finally, column “49. iteration” displays the final $u_{i}$-values. 9286 out of 9600 of the contaminated training observations have achieved a $u_{i}\in(-2.08,-1.56]\cdot 10^{-5}$. This means these training observations are removed, or nearly-so, from consideration because this value cancels the nominal probability $1/N$ = 2.08 $\cdot$ $10^{-5}$. It is observed that a large majority (35246/38400) clean training observations remain with their nominal probability. This helps explain the performance benefit of A-RRM over AT. A-RRM ”removes” the contaminated data points in-situ, whereas AT does not. It appears that under adversarial training regimes with contaminated training data, it is essential to identify and ”remove” the contaminated examples, especially if the level adversarial perturbation encountered in the test set is unknown, or possibly lower than the level of adversarial perturbation applied to the training set.
Table 5: MNIST-10 Evolution of $u$ across $|C|=9600$ contaminated data points and $N-|C|=38400$ clean data points. Note that $1/N=2.08\cdot 10^{-5}$.
1. iteration 10. iteration 49. iteration $C$ $[N]\setminus C$ $C$ $[N]\setminus C$ $C$ $[N]\setminus C$ $u_{i}$ value ($10^{-5}$) $\gg$ 0 0 1 0 4 0 25 $\approx$ 0 8844 38385 2058 37524 91 35246 (-0.52, 0.00) 0 0 7 36 146 1655 (-1.04, -0.52] 0 0 41 45 43 155 (-1.56, -1.04] 756 14 415 174 34 168 (-2.08, -1.56] 0 0 7079 617 9286 1151
@@ 7 Conclusion
A-RRM/RRM algorithm exhibits robustness in a variety of data domains, data contamination schemes, model architectures and machine learning applications. In truth, we have presented a meta-algorithm that operates on loss function methodologies, and further has the capacity to incorporate benchmarks/estimates of label contamination and adversarial feature perturbation. Furthermore, the method is virtually hyperparameter-less, upon utilizing an auto-tuning feature that enables control over the data-pruning action of Re-weight, to be used with an estimate of the label contamination level.
Our experiments indicate RRM not only amplifies test accuracy of methods that don’t account for contamination, but even those already robust to contamination. In the MNIST-10 example we show that conducting training in preparation for deployment environments with varied levels of adversarial attacks, one can benefit from implementation of the A-RRM algorithm. This can lead to a model more robust across levels of both feature perturbation and high levels of label contamination. We also demonstrate the mechanism by which A-RRM operates and confers superior results: by automatically identifying and removing the contaminated training observations at training time execution.
Impact Statement
This work seeks to advance the field of Machine Learning, particularly in the area of model robustness. While various secondary implications may emerge from its application, we do not identify any immediate or critical concerns that warrant specific emphasis in this context.
@@ A.1 Section 3 Proofs
$h(u)=\sum_{i=1}^{N}\Big{(}c_{i}/N+u_{i}c_{i}+\frac{\gamma}{2}|u_{i}|+\iota_{[0,\infty)}(1/N+u_{i})\Big{)}+\iota_{\{0\}}\Big{(}\sum_{i=1}^{N}u_{i}\Big{)}$
Thus, because $h(u)>-\infty$ for all $u\in\mathbb{R}^{N}$ and $h$ is convex, $u^{\star}$ is a solution of the minimization problem if and only if $0\in\partial h(u^{\star})$. We proceed by characterizing $\partial h$.
@@
$h_{i}(u_{i})=c_{i}/N+u_{i}c_{i}+\frac{\gamma}{2}|u_{i}|+\iota_{[0,\infty)}(1/N+u_{i}).$
For $u_{i}\geq-1/N$, the Moreau-Rockafellar sum rule reveals that
$\partial h_{i}(u_{i})=c_{i}+\begin{cases}\{\frac{\gamma}{2}\}&\mbox{ if }u_{i}>0\\ [-\frac{\gamma}{2},\frac{\gamma}{2}]&\mbox{ if }u_{i}=0\\ \{-\frac{\gamma}{2}\}&\mbox{ if }-1/N<u_{i}<0\\ (-\infty,-\frac{\gamma}{2}]&\mbox{ if }u_{i}=-1/N.\end{cases}$
For $u=(u_{1},\dots,u_{N})\in[-1/N,\infty)^{N}$, we obtain by Proposition 4.63 in (Royset & Wets, 2021) that
@@
$\lambda\in\begin{cases}\{c_{i}+\frac{\gamma}{2}\}&\mbox{ if }u_{i}^{\star}>0\\ [c_{i}-\frac{\gamma}{2},c_{i}+\frac{\gamma}{2}]&\mbox{ if }u_{i}^{\star}=0\\ \{c_{i}-\frac{\gamma}{2}\}&\mbox{ if }u_{i}^{\star}\in(-1/N,0)\\ (-\infty,c_{i}-\frac{\gamma}{2}]&\mbox{ if }u_{i}^{\star}=-1/N.\end{cases}$ (6)
It follows that any such $\lambda$ must then satisfy $\lambda\leq c_{\min}+\frac{\gamma}{2}$, as the above list of cases indicate for arbitrary $i$ and no matter the value of $u_{i}^{*}$, for any optimal $u^{*}\in U$.
We proceed to show that in fact, $u^{*}\in U$ solves (4) if and only if $\lambda=c_{min}+\frac{\gamma}{2}$ satisfies the conditions of (6) with respect to $u^{*}$ for all $i$. Suppose $u^{*}\in U$ is optimal and that there exists a corresponding $\lambda$ satisfying (6) for all $i$, and that it satisfies $\lambda<c_{min}+\frac{\gamma}{2}.$ Then for any optimal $u^{*}\in U$, there exists no $i$ for which $u_{i}^{*}>0,$ otherwise $c_{i}=\lambda-\frac{\gamma}{2}<c_{min},$ an impossibility. This means any optimal $u^{*}\in U$ must satisfy $u^{*}\leq 0,$ for which then we would conclude that in fact $u^{*}=0,$ and is the unique optimal solution, which in turn yields that $\lambda-\frac{\gamma}{2}\leq c_{i}\leq\lambda+\frac{\gamma}{2}<c_{min}+\gamma$ for all $i.$ This last inequality reveals that $\lambda^{*}:=c_{min}+\gamma$ is compatible with the unique optimal solution $u^{*}=0\in U$ in such a case. In summary, if a $\lambda<c_{min}+\frac{\gamma}{2}$ is compatible with all optimal solutions $u^{*}\in U,$ then $\lambda^{*}:=c_{min}+\frac{\gamma}{2}$ is as well.
It follows that $u^{*}\in U$ solves (4) if and only if $\lambda=c_{min}+\frac{\gamma}{2}$ satisfies the conditions of (6) with respect to $u_{i}^{*}$ for all $i$; hence, the result follows. ∎
See 3.1.1
Proof.
Clearly, $u^{*}\in U$. Let $c_{i}=J(\theta^{*};x_{i},y_{i})$ for $i=1,\ldots,N.$ In the proof of Theorem 3.1, it was shown that $u\in U$ solves (4) if and only if $\lambda=c_{min}+\frac{\gamma}{2}$ satisfies the conditions of (6) with respect to $u_{i}*$ for all $i$. Therefore, with $S_{1}=S_{2}=\emptyset,$ Theorem 3.1 indicates that $u^{*}\in U^{*}_{S_{1},S)_{2}}$ and hence must solve (4). ∎
See 3.1
@@
By Lemma 3.1 of (Staib & Jegelka, 2017) and/or Corollary 2 of (Gao & Kleywegt, 2023),
@@
Appendix B Non-uniform label Contamination Experiment
Although we explicitly state the use of uniform label noise in Section 3.1, which is indeed a very common scheme in the literature, our analysis in fact did not rely on this assumption. Towards providing insight into the non-uniform case, we have repeated the experiments of Section 6.3 that produced Table 4, but now with non-uniform label noise. More precisely, after uniformly at random selecting $C$ percent of the training pairs, we proceed to contaminate the label $y_{i}$ in each pair $(x_{i},y_{i})$ in the following non-uniform manner, as outlined below in the transition kernel matrix of (True Label, Contaminated Label) entries. For example, if the true label $y_{i}=5$, then instead of uniformly at random drawing an alternative digit $\tilde{y}_{i}$ from among $\{0,1,\ldots,9\}\setminus\{5\}$ we have $\tilde{y}_{i}=\begin{cases}0&w.p.0.051\\ 1&w.p.0.017\\ 2&w.p.0.\\ 3&w.p.0.627\end{cases}$
Table 6: Contamination Kernel
Contaminated
0 1 2 3 4 5 6 7 8 9
Original 0 0 0.0769 0.0769 0.1538 0 0.0769 0.3846 0 0.1538 0.0769
1 0 0 0.3333 0.1111 0 0.1111 0.1111 0 0.3333 0
2 0.0968 0.0645 0 0.2581 0.0323 0 0.0968 0.1935 0.2581 0
3 0 0 0.1250 0 0 0.1250 0 0.1250 0.6250 0
4 0.1111 0.0370 0.0741 0.0741 0 0.0741 0.2222 0.0370 0.1111 0.2593
5 0.0508 0.0169 0 0.6271 0.0169 0 0.1525 0 0.1017 0.0339
6 0.2353 0.1765 0.0588 0.0588 0.0588 0.1765 0 0 0.2353 0
7 0.0500 0.2250 0.2000 0.1250 0 0 0 0 0.2000 0.2000
8 0.1071 0.0357 0.1071 0.3571 0.1071 0.0714 0.0714 0.1071 0 0.0357
9 0.0638 0.1702 0 0.2128 0.1702 0.1702 0.0213 0.0851 0.1064 0
These entries were generated by the confusion matrix of an imperfect MNIST classifier. The results from this new experiment confirm the performance benefits that were observed (compare to Table 4 under conditions of uniform label contamination.
Table 7: MNIST-10 Test accuracy (%) for AT and A-RRM under different levels of contamination $C$, training perturbation $\epsilon=1.0$, and test-set adversarial perturbation $\epsilon_{test}$.
$C$
0% 5% 10% 20% 30%
AT A-RRM AT A-RRM AT A-RRM AT A-RRM AT A-RRM
$\epsilon_{test}$ 0.00 96.5 97.3 93.6 95.6 60.4 87.8 32.2 92.4 58.3 89.2
0.10 93.4 95.2 89.3 92.4 63.2 84.7 42.5 89.5 56.1 81.7
0.25 92.4 93.1 87.9 90.6 86.9 86.3 80.6 89.3 69.0 79.8
0.50 92.0 90.9 90.3 89.8 94.4 91.2 92.9 89.2 85.2 81.6
1.00 89.4 85.5 90.3 86.8 94.9 92.6 93.9 86.6 81.8 77.8
Appendix C Additional Data Experiments
C.1 Toxic Comments
Dataset: Toxic Comments [^3] https://kaggle.com/competitions/jigsaw-toxic-comment-classification-challenge is a multi-label classification problem from JIGSAW that consists of Wikipedia comments labeled by humans for toxic behavior. Comments can be any number (including zero) of six categories: toxic, severe toxic, obscene, threat, insult, and identity hate. We convert this into a binary classification problem by treating the label as either none of the six categories or at least one of the six categories. This dataset is a public dataset used as part of the Kaggle Toxic Comment Classification Challenge.
Architecture: We use a simple model with only a single convolutional layer. A pretrained embedding from FastText is first used to map the comments into a 300 dimension embedding space, followed by a single convolutional layer with a kernel size of two with a ReLU activation layer followed by a max-pooling layer. We then apply a 36-unit dense layer, followed by a 6 unit dense layer with sigmoid activation. Binary cross-entropy is used for the loss function.
Experiment: We use the Toxic Comments dataset to test the efficacy of RRM on low prevalence text data. The positive (toxic) comments consist of only 3% of the data and we contaminate anywhere from 1% to 20% of the labels. There are a total of 148,000 samples, and we set aside 80% for training and 20% for test. $\sigma=2$ with 3 iterations of the heuristic algorithm results in a total of 6 epochs, and ERM is run for a total of 6 epochs to make the results comparable. Since the data is highly imbalanced, we look at the area under the curve of the precision/recall curve to assess the performance of the models. Unsurprisingly, as the noise increase, the model performance decreases. We note that RRM outperforms ERM across all noise levels tested, though as the noise increase, the gap between RRM and ERM decreases.
Table 8: Toxic Comments Comparison of training and test area under the precision/recall curve for ERM and RRM at noise levels ranging from 1% to 20%.
Method Percentage Contaminated Training Data 1% 5% 7% 10% 15% 20% ERM (train) 0.2904 0.2006 0.1589 0.1302 0.1073 0.0920 RRM (train) 0.6875 0.4458 0.3805 0.3087 0.2438 0.1966 ERM (test) 0.5861 0.3970 0.3246 0.2550 0.2013 0.1717 RRM (test) 0.6705 0.4338 0.3619 0.2824 0.2208 0.1861
Takeaway: The Toxic Comment example presents another challenging classification problem, characterized by a low prevalence target class amidst label noise. Our experiments demonstrate that as the amount of label noise increases, standard methods become increasingly ineffective. However, RRM remains reasonably robust under varying degrees of label contamination. Therefore, RRM could be a valuable addition to the set of tools being developed to enhance the robustness of AI-based decision engines.
C.2 IMDb
Dataset (Maas et al., 2011): A binary classification dataset consisting of 50000 movie reviews each assigned a positive or negative sentiment label. 25000 reviews are selected randomly for training and the remaining are used for testing. 25%, 30%, 40%, and 45% of the labels of the training reviews are randomly selected and swapped from positive sentiment to negative sentiment, and vice versa, to achieve four training datasets of desired levels of label contamination.
Architecture: Transformer architectures have achieved SOA performance on the IMDb dataset sentiment analysis task (Devlin et al., 2019; Xie et al., 2020). As such, we a adopt a reasonable transformer architecture to assess RRM. We utilize the DistilBERT (Sanh et al., 2020) architecture with low-rank adaptation (LoRA) (Hu et al., 2022) for large language models, which reduces the number of trainable weights from 67584004 to 628994. In this manner, we reduce the computational burden, while maintaining excellent sentiment analysis performance. Binary cross-entropy is employed for the loss function.
Experiment: Twenty percent of the training data is set aside for validation purposes. Using Pytorch 2.1.0, 30 iterations of RRM are executed, with $\sigma=10$ epochs per iteration for a total of 300 epochs for a given hyperparameter setting. For RRM, the hyperparameter settings of $\mu$ and $\gamma$ at 0.5 and 0.4, respectively, are based on a search to optimize validation set accuracy. For contrast, we perform a comparable 300 epochs using ERM. Both ERM and RRM employ stochastic gradient descent (SGD) with a learning rate ($\eta$) of $0.001$. In Table 9 we record both the test set accuracy achieved when validation set accuracy peaks, as well as the maximum test set accuracy. At these high levels of contamination RRM consistently achieves a better maximum test set accuracy.
Table 9: IMDb Test accuracy (%) for ERM and RRM under different levels of contamination. Test set accuracy at peak validation accuracy and maximum test set accuracy are recorded.
Method Percentage Contaminated Training Data
25% 30% 40% 45%
ERM 90.2, 90.2 89.5, 89.6 86.4, 86.6 80.7, 81.1
RRM 90.1, 90.4 90.2, 90.4 88.4, 88.7 76.9, 82.6
Takeaway: We demonstrate that RRM can confer benefits to the sentiment analysis classification task using pre-trained large models under conditions of high label contamination. The success of fine-tuning in LLMs depends, in large part, on access to high quality training examples. We have shown that RRM can mitigate this need by allowing effective training in scenarios of high training data contamination. As such, resource allocation dedicated to dataset curation may be lessened by the usage of RRM.
C.3 Tissue Necrosis
Dataset: A binary classification dataset consisting of 7874 256x256-pixel hematoxylin and eosin (H&E) stained RGB images derived from (Amgad et al., 2019). The training dataset consists of 3156 images labeled non-necrotic, as well as 3156 images labeled necrotic. The training images labeled non-necrotic contain no necrosis. However, only 25% of the images labeled necrotic contain necrotic tissue. This type of label error can be expected in cases of weakly-labeled Whole Slide Imagery (WSI). Here, an expert pathologist will provide a slide-level label for a potentially massive slide consisting of gigapixels, but they lack time or resources to provide granular, segmentation-level annotations of the location of the pathology in question. Also, the diseased tissue often occupies a small portion of the WSI, with the remainder consisting of normal tissue. When the gigapixel-sized WSI is subsequently divided into sub-images of manageable size for typical machine-learning workflows, many of the sub-images will contain no disease, but will be assigned the ”weak” label chosen by the expert for the WSI. The test dataset consists of 718 necrosis and 781 non-necrosis 256x256-pixel H&E images, which were also derived from (Amgad et al., 2019). For both the training and test images, (Amgad et al., 2019) provide segmentation-level necrosis annotations, so we are able to ensure a pristine test set, and, in the case of the training set, we were able to identify the contaminated images for the purpose of algorithm evaluation.
Architecture: Consistent with the computational histopathology literature (Petríková & Cimrák, 2023), we employ a convolutional neural network (CNN) architecture for this classification task. In particular, a ResNet-50 architecture with pre-trained ImageNet weights is harnessed. The classification head is removed and replaced with a dense layer of 512 units and ReLU activation function, followed by an output layer with a single unit using a sigmoid activation function. All weights, with the exception of the new classification head are frozen, resulting in 1050114 trainable parameters out of 24637826. Binary cross-entropy is employed for the loss function.
Experiment:Twenty percent of the training data is set aside for validation purposes, including hyperparameter selection. 60 iterations of RRM are executed, with $\sigma=10$ epochs per iteration, for a total of 600 epochs for a given hyperparameter setting. For RRM, the hyperparameter settings of $\mu$ and $\gamma$ at 0.5 and 0.016, respectively, are based on a search to optimize validation set accuracy. For contrast, we perform a comparable 600 epochs using ERM. Both ERM and RRM employ stochastic gradient descent (SGD) with a learning rate ($\eta$) of $5.0$ and $1.0$, respectively. RRM achieves a test set accuracy at peak validation accuracy of 74.6, and a maximum test set accuracy 77.2, whereas ERM achieves 71.7 and 73.2, respectively. RRM appears to confer a performance benefit under this regime of weakly labeled data.
Takeaway: In the Tissue Necrosis example, we demonstrate that RRM also confers accuracy benefits to the necrosis identification task provided weakly labeled WSIs. Again, RRM can mitigate the need for expert-curated, detailed pathology annotations, which are costly and time-consuming to generate.